Calculate liters of CO2 from grams of octane
1
Expert's answer
2018-01-16T14:14:06-0500
Mr (C8H18) = 114g/mol
1 g x L
2C8H18 + O2 = 16CO2 + 18H2O
2 * 114 g 16 * 22.4 L

x = 1 * 16 * 22.4 / 2 / 114 = 1.57 L

Answer:

1.57 L of CO2 forms from 1 g of octane